# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ 3 Constant number of players
If the number of players is bounded by a constant $m$, then the emptiness of the weak (resp. strong) core can be decided in polynomial time. Furthermore, if the weak (resp. strong) core is nonempty, then a weak (resp. strong) core solution can also be found in polynomial time.
@@ 4 Hardness results for partition classes of constant size
In contrast to the case where the number of players is bounded by a constant, bounding the sizes of the partition classes by a constant does not lead to polynomial-time solvable problems. In this section, we show hardness results of this type. We have seen in Section 2 that our problems are tractable when all partition classes have size at most 2. Here we show that having partition classes of size 3 leads to coNP-hardness for most problems.
@@ 4.1 Weak core
In this section we focus on the weak core. Our first result is the hardness of verifying if a matching belongs to the weak core.
@@
It is coNP-complete to decide if a matching $M$ is in the weak core, even if each partition class has size at most 3 and the graph is bipartite.
@@
Let $I$ be an instance of x3c. We create a bipartite instance $I^{\prime}$ of the NTU international matching game as follows. First, we describe the vertex set $V$ of the graph.
  • For each element $i\in[3n]$ we create an element vertex $a_{i}$.
@@
Finally, we describe the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2}\}$. For each $i\in\{2,\dots,3n-1\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$. In addition, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, and $A_{3n}=\{b_{3n-1,3n}^{3n},a_{3n}\}$. Clearly, all players have size at most 3.
We create a matching $M$ as follows. We add $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[2]$ and $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i\in[3n-1]$.
@@
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $\{C_{l_{1}},\dots,C_{l_{n}},A_{1},\dots,A_{3n}\}$ form a strongly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $(b_{i,i+1}^{i},b_{i,i+1}^{i+1})$ for $i=1,\dots,3n-1$. Hence, there is a matching $M^{\prime}$ among these players that matches every single one of their vertices. As each of these players had a vertex exposed in $M$ ($c_{j}^{3}$ for $C_{j}$, $a_{i}$ for $A_{i}$), it follows that $M$ is not in the weak core. ∎
@@
Suppose that $M$ is not in the weak core. This means that there is a strongly blocking coalition $\mathcal{P}$ to $M$ with a matching $M^{\prime}$. We can suppose that there is a player $A_{i}\in\mathcal{P}$. Otherwise, there could be only $C_{j}$ players in $\mathcal{P}$ (the $D_{j}$ players have all vertices covered by $M$), and they induce no edges between them, contradiction.
Observe that players $A_{1},\dots,A_{3n}$ are all in $\mathcal{P}$, because for $A_{i}$ to strictly improve, $b_{i-1,i}^{i}$ and $b_{i,i+1}^{i}$ (if they exists) must be matched, so $b_{i-1,i}^{i-1}b_{i-1,i}^{i},b_{i,i+1}^{i}b_{i,i+1}^{i+1}\in M^{\prime}$. This implies that $A_{i-1},A_{i+1}\in\mathcal{P}$ and similarly we get that $A_{1},\dots,A_{3n}$ are all in $\mathcal{P}$.
Finally, as each player must strictly improve, every $a_{i}$ vertex is matched in $M^{\prime}$ and so they are matched to $C_{j}$ players. Whenever a $C_{j}$ player participates in the blocking by matching one of its vertices to an $a_{i}$ vertex, $\{c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j_{3}}\}\subseteq M^{\prime}$, because $D_{j}\notin\mathcal{P}$ and $C_{j}$ must strictly improve. Combining this with that fact that each $a_{i}$ is matched, we obtain that these players must correspond to an exact 3-cover. ∎
@@
Intuitively, it would be tempting to conjecture that the weak core is always nonempty in any NTU partitioned matching game. However, the following example (given by Zsuzsanna Jankó) shows that this is not the case.
Example 20.

### v2

@@ 2.2.1 Deciding the existence of a strong core matching
For the subsets of players $\mathcal{K}^{*}$ and $\mathcal{L}$, we have that $\mathcal{K}^{*}\subseteq\mathcal{L}$.
Lemma 15.
Let $A,B,C\in\mathcal{K}$ and consider an $(A,B;C)$-$\delta$-path $H$, and an $M\in\mathcal{SC}$ such that $|V(M)\cap A|=1$ and $|V(M)\cap B|=2$. Then $e_{A}$ and $e_{B}$ are on the same alternating path in $M\Delta M_{0}$.
Proof.
Suppose that $e_{A}$ is on the $M\Delta M_{0}$-alternating path $P_{A}$ and $e_{B}$ is on $M\Delta M_{0}$-alternating path $P_{B}$ and $P_{A}\neq P_{B}$. By $|V(M)\cap A|=1$, $a\in A$ is on one end of $P_{A}$; let the other end be $d\in D$. Let the endpoints of $P_{B}$ be $x_{B}\in X$ and $y_{B}\in Y$. Using Lemma 11, we get that we can assume that $|V(M)\cap C|=1$.
Let $P_{BC}$ be the $M_{0}$-alternating path of $H$ starting with $e_{B}$ and ending with $e_{C}$ that contains $e_{A}$. We use Lemma 9 as follows. We set $A_{L}:=A,B_{L}=C_{L}:=D$, $P^{L}$ as $P_{A}$, and $P^{L}_{A_{L}}$ as taking a maximal subpath of $P_{BC}$ that contains $e_{A}$, but does not intersect $P_{B}$ (as $e_{A}\in P_{BC}$, but $e_{A}\notin P_{B}$, there must be such a subpath). Then, Lemma 9 gives a path $P_{A}^{\prime}$ disjoint from $P_{B}$ (neither $P^{L}$ nor $P^{L}_{A_{L}}$ intersects $P_{B}$) that ends in $d\in D$, contains $e_{A}$ and either the other end is connected by an edge to $P_{B}$, or if not, then it ends in $C$ (as $e_{B}\in P_{B}$). In both cases, we obtain an $M_{0}$-alternating path that contains $e_{A}$ in its interior, and connects two players who both have one vertex covered by $M$. Thus, the players on this path form a weakly blocking coalition, contradicting $M\in\mathcal{SC}$. ∎
Lemma 16.
Proof.
Suppose for the contrary that $(B,C)$ is not a pair. Let $M\in\mathcal{SC}$ be a matching such that $|V(M)\cap A|=1$, and let $P$ be the path in $M\Delta M_{0}$ that contains $e_{A}$. By Lemma 15, $e_{B}$ and $e_{C}$ are on $P$ and $|V(M)\cap B|=|V(M)\cap C|=2$. We may assume that $P$ is an ABCDA-B-C-D path, where $D\in\mathcal{K}^{*}$ and $|V(M)\cap D|=1$. Let the order of the vertices of these players along the path be $a,a^{\prime},b,b^{\prime},c,c^{\prime},d,d^{\prime}$.
As $(B,C)$ is not a pair, but $C\in\mathcal{L}$, we get that there must be a $(C,D;X)$-$\delta$-path $H$, by using the definition of $\mathcal{L}$ for the BCDB-C-D subpath of $P$. We may assume that $|V(M)\cap X|\leq 1$ and $X\in\mathcal{K}^{*}$ by Lemma 11.
We use Lemma 9 as follows. We set $A_{L}:=D$, $B_{L}:=C$ and $C_{L}:=A$. We let $P^{L}:=P$ and $P^{L}_{A_{L}}$ be the CDXC-D-X $M_{0}$-alternating path of the $\delta$-path $H$. If we get an $M_{0}$-alternating path through $e_{D}$ that ends in $A$ and starts in $X$, then that corresponds to a blocking coalition to $M$, contradiction. Hence, we get a $(D,C;A)$-$\delta$-path $H^{\prime}$, such that it contains the ABCA-B-C subpath of $P$.
As $(A,C)$ is a pair, we know that there is an $(A,C;Y)$-$\delta$-path $H^{\prime\prime}$ for some $Y\in\mathcal{K}^{*}$ with $|V(M)\cap Y|\leq 1$ by Lemma 11.
Finally, we use Lemma 9 for $A_{L}:=A$, $B_{L}:=Z$ and $C_{L}:=D$, where $Z$ is the unique player with the connector vertex in the $(D,C;A)$-$\delta$-path $H^{\prime}$. We set $P^{L}$ to be the ABCDA-B-C-D subpath of $H^{\prime}$ and $P^{L}_{A_{L}}$ to be the maximal subpath of the CAYC-A-Y $M_{0}$-alternating path of $H^{\prime\prime}$, such that it contains $e_{A}$, but does not intersect the cycle $\mathrm{cyc}(H^{\prime})$ of $H^{\prime}$. As $e_{A}\notin\mathrm{cyc}(H^{\prime})$, there is such a subpath. Then, Lemma 9 gives one of two options. In the first one, we get an YADY-A-D $M_{0}$-alternating path, where all three of these players have at most 1 vertex covered by $M$, contradicting $M\in\mathcal{SC}$. In the second option, we get $u\in A$, and path $P_{uz}$, such that the end of $P_{u}^{L}$ is connected with an edge to $\mathrm{cyc}(H^{\prime})$ (since the first option does not hold, and $e_{C}\in P^{L}$). However, this implies that we obtain an $M_{0}$-alternating cycle through $e_{A}$, contradiction. ∎
Corollary 17.
We have that $\mathcal{K}^{*}\subseteq\mathcal{L}^{*}$.
Proof of Theorem 8.
To see that the conditions are necessary, observe that $|V(M)\cap A|=2$ must hold for every $A\notin\mathcal{K}^{*}$, hence also for $A\notin\mathcal{L}^{*}$ as $\mathcal{K}^{*}\subseteq\mathcal{L}^{*}$. Also, if $(A,B)$ is a pair, then $|V(M)\cap(A\cup B)|\geq 3$ must hold by Corollary 12. Lastly, if $A\notin\mathcal{K}_{0}$, then there is a player $B\in\mathcal{K}$ such that there is an ABA-B $M_{0}$-alternating path, hence $|V(M)\cap A|\geq 1$ must hold by Lemma 10.
To see sufficiency, we use the alternating path characterization in Lemma 6 for membership in $\mathcal{SC}$. Suppose $M$ satisfies the two conditions of Theorem 8. Then, the properties of $M$ imply that if a player-edge is in an $M_{0}$-alternating cycle, then both of its vertices are covered by $M$. Let $P$ be an $M_{0}$-alternating path from $A\in\mathcal{K}$ to $B\in\mathcal{K}$, where none of the vertices of $A$ are covered by $M$. Then $A\in\mathcal{K}_{0}$, which contradicts that there is an $M_{0}$-alternating path to some $B\in\mathcal{K}$ from $A$. Finally, suppose that $P$ is an ABCA-B-C alternating path, where $|V(M)\cap B|=1$. If $A\notin\mathcal{L}^{*}$, $B\notin\mathcal{L}^{*}$ or $C\notin\mathcal{L}^{*}$, then one of $A,B,C$ is covered twice by $M$, and we are done. Otherwise, since $A,C\in\mathcal{L}^{*}\subseteq\mathcal{K}$ and $B\in\mathcal{L}^{*}\subseteq\mathcal{L}$, the definition of $\mathcal{L}$ implies that $(A,B)$ is a pair or $(B,C)$ is a pair. In the former case, $|V(M)\cap A|=2$, while the latter case implies $|V(M)\cap C|=2$ by the second condition on $M$. Thus, the conditions of Lemma 6 are satisfied and therefore $M$ is in the strong core. ∎
To obtain polynomial-time algorithms, it remains to show that membership in $\mathcal{L}$ and $\mathcal{L}^{*}$ can be decided in polynomial time. Let $\alpha$ denote the inverse Ackermann function. Then, it is known that detecting a negative weight cycle in an undirected graph with integer edge costs bounded by a constant can be done in $\mathcal{O}(|V|^{2.5}\log(|V|)^{1.5}\alpha(|V|^{2},|V|)^{0.5})$ time [50] (Corollary 2). Hence, for any constant $\varepsilon>0$ it is true that it can be done in $\mathcal{O}(|V|^{2.5+\varepsilon})$ time. We will use this fact in our next lemma.
Lemma 18.
For given $A,B,C\in\mathcal{K}$, it can be decided in $O(|V|^{2.5+\varepsilon})$ time if an $(A,B;C)$-$\delta$-path exists.
Proof.
As mentioned in the proof of Theorem 7, we can decide in $\mathcal{O}(|V|^{2.5})$ time if there is an ABCA-B-C $M_{0}$-alternating path for $A,B,C\in\mathcal{K}$. To make sure that $B$ is the middle player, we can iterate through all pairs $(u,v)\in\{a,a^{\prime}\}\times\{c,c^{\prime}\}$ and delete the edges in $E$ that are adjacent to $u$ and $v$, when searching for an almost-perfect matching.
If there is no such $M_{0}$-alternating path, then we can conclude that no $(A,B;C)$-$\delta$-path exist. Otherwise, let $P$ be the ABCA-B-C $M_{0}$-alternating path found, and suppose the order of the vertices is $a,a^{\prime},b,b^{\prime},c,c^{\prime}$.
Suppose that there exists an $(A,B;C)$-$\delta$-path $H$. Using Lemma 9 for $A_{L}:=A,B_{L}:=B$ and $C_{L}:=C$, $P^{L}:=P$ and the BACB-A-C path subpath of $H$ as $P^{L}_{A_{L}}$, we obtain that there must exist an $(A,B;C)$-$\delta$-path $H^{\prime}$ that contains the part of $P$ from $e_{B}$ to $e_{C}$. Thus, it is enough to decide if there exists such an $(A,B;C)$-$\delta$-path or not. To do this, first, we create a new graph from $G$ as follows.
  1. 1.
    Add $M_{0}$.
  2. 2.
    Delete $c^{\prime}$.
  3. 3.
    Contract all vertices strictly after $b^{\prime}$ on $P$ into $c$. Let this new vertex be $\hat{c}$.
  4. 4.
    Delete all incident edges to $b^{\prime}$, except $e_{B}$ and $b^{\prime}\hat{c}$.
Let the graph obtained be $\hat{G}$. We assign a weight of 1-1 to the edges of $M_{0}$ in $\hat{G}$, except that we assign 2-2 to $e_{A}$ and $e_{B}$. All other edges have a weight of $1$.
We claim that there is a negative weight cycle in $\hat{G}$ if and only if there is an $(A,B;C)$-$\delta$-path in $E\cup M_{0}$. Since the former can be decided in polynomial time, the statement will also follow from this.
Suppose that there is an $(A,B;C)$-$\delta$-path $H$; we can assume that the part $P_{bc^{\prime}}$ of $P$ from $b$ to $c^{\prime}$ is included in $H$. It is easy to see that $(H\setminus P_{bc^{\prime}})\cup\{e_{B},b^{\prime}\hat{c}\}$ will give a cycle in $\hat{G}$ with weight 1-1, as it alternates between $E$ and $M_{0}$ except from $\hat{c}$ and contains both $e_{A}$ and $e_{B}$.
Conversely, suppose that there is a negative weight cycle in $\hat{G}$. As there is no player edge adjacent to $\hat{c}$, no $M_{0}$-alternating cycle goes through $\hat{c}$. Hence, as $A,B\in\mathcal{K}$, there cannot be $M_{0}$-alternating cycles in $\hat{G}$ that contain either $e_{A}$ or $e_{B}$. Since $M_{0}$ forms a matching, any negative weight cycle must contain $e_{A}$ or $e_{B}$. Furthermore, any negative weight cycle that is not $M_{0}$-alternating must contain both. Hence, by the previous observation, there must be a cycle in $\hat{G}$ that is $M_{0}$-alternating except at one vertex $v$, and contains $e_{A}$ and $e_{B}$. Since $b^{\prime}$ has only $e_{B}$ and $b^{\prime}\hat{c}$ as incident edges, we must have that $v=\hat{c}$. Then, by $A,B\in\mathcal{K}$, it is easy to see that this cycle gives an $(A,B;C)$-$\delta$-path in $E\cup M_{0}$ (by adding back $e_{C}$ and undoing the contractions).
As finding a negative cycle in an undirected graph with edge weights $\{-2,-1,1\}$ can be done in $O(|V|^{2.5+\varepsilon})$ time [50], the lemma follows. ∎
The bottleneck for the running time of the algorithms is the verification of membership in $\mathcal{L}$ and $\mathcal{L}^{*}$, and the construction of the graph $G^{*}$. We did not try to optimize the running time of these subroutines; the naive approach requires $O(|V|^{3})$ maximum size matching algorithms and negative cycle detections in an undirected graph with edge costs in $\{-2,-1,1\}$. Since both of these can be done in time $O(|V|^{2.5+\varepsilon})$ [19, 50], we get a total running time of $O(|V|^{5.5+\varepsilon})$.
Corollary 19.
NTU-PMG-SC-Exist-2 is solvable in $O(|V|^{5.5+\varepsilon})$ time.
Optimization on the strong core.
An interesting and useful consequence of Theorem 8 is that for given edge costs $c\in\mathbb{Q}^{E}$, we can find a minimum cost matching in the strong core. This is in stark contrast to the case of the weak core, where optimization is NP-hard even for players of size 1, as shown in Theorem 5. Indeed, by Theorem 8, the strong core can be characterized by setting lower bounds on $|U_{i}\cap V(M)|$ for some partition $U_{1}\cup\cdots\cup U_{l}$, so by Lemma 1, after the partition and the bounds are computed, a minimum cost such matching can be found in $O(|V|^{4})$ time, if one exists.
Corollary 20.
If the strong core is nonempty, then for any edge costs $c\in\mathbb{Q}^{E}$, we can find a minimum-cost matching in the strong core in $O(|V|^{5.5+\varepsilon})$ time.
3 Small number of players
In this section, we show that if the number of players is small, i.e., bounded by a constant, then the existence of a strong (resp. weak) core solution, the verification of membership in the cores, and optimization on the cores are all solvable in polynomial time, with an algorithm that is only exponential in the number of players. This is relevant for example in the IKEP application, where players correspond to countries, which often implies a fairly small upper bound on the number of players, usually between 3 and 8 [14, 49]. We shall see that the tractability of this case is a direct consequence of Lemma 1.
Recall that the instance is given by a graph $G=(V,E)$ and a partition $\Pi=(V_{1},V_{2},\dots,V_{m})$ of $V$, where $m$ is the number of players. In this section, $m$ is regarded as a small constant, but our results are true for arbitrary $m$. However, the running times become infeasible for large $m$, because of the exponential dependence on $m$.
First, we consider the verification problems. Lemma 1 immediately implies that membership of a given matching $M$ in the weak core can be decided in $\mathcal{O}(2^{m}|V|^{2.5})$ time. Indeed, the number of possible coalitions is $2^{m}$, and for a given coalition $\{i_{1},i_{2},\dots,i_{k}\}$, we can decide, using Lemma 1, if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|+1$ for every $j\in[k]$ in $\mathcal{O}(|V|^{2.5})$ time.
Membership in the strong core can be decided similarly, but for every coalition $\{i_{1},i_{2},\dots,i_{k}\}$ and every $\ell\in[k]$, we check if there is a matching $M^{\prime}$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M^{\prime})\cap V_{i_{j}}|\geq|V(M)\cap V_{i_{j}}|$ for every $j\in[k]$ and $|V(M^{\prime})\cap V_{i_{\ell}}|\geq|V(M)\cap V_{i_{\ell}}|+1$. This leads to a runtime of $\mathcal{O}(2^{m}|V|^{3.5})$.
Corollary 21.
NTU-PMG-WC-Verif can be solved in $\mathcal{O}(2^{m}|V|^{2.5})$ time, and NTU-PMG-SC-Verif can be solved in $\mathcal{O}(2^{m}|V|^{3.5})$ time.
Deciding non-emptiness of the weak and strong cores is somewhat more difficult, but still polynomial-time solvable if $m$ is constant. The crucial observation is that the membership of a matching $M$ in these cores depends only on the values $|V(M)\cap V_{i}|$ ($i\in[m]$). Let us call an integer vector maximal with respect to a property if the property is violated by increasing any component of the vector by 1. Using Lemma 1, we can find in $\mathcal{O}(|V|^{m+2.5})$ time all maximal vectors $x\in\mathbb{Z}_{+}^{m}$ for which a matching $M$ with $|V(M)\cap V_{i}|\geq x_{i}$ for every $i\in[m]$ exists. Indeed, $x_{i}\leq|V_{i}|$ ($i\in[m]$) can be assumed, so we can check all possible $\mathcal{O}(|V|^{m})$ vectors, each in $\mathcal{O}(|V|^{2.5})$ time. Let $\mathcal{X}$ denote the set of such maximal vectors. Since for any matching $M$ covering a set $U\subseteq V$ of vertices, there exists a maximum size matching $M^{\prime}$ that also covers $U$, and so $|V(M^{\prime})\cap V_{i}|\geq|V(M)\cap V_{i}|$ $\forall i\in[m]$, we can assume that for each $x\in\mathcal{X}$, the sum of its coordinates is two times the size of a maximum matching. Hence, $|\mathcal{X}|\leq\mathcal{O}(|V|^{m-1})$
It is easy to see that if the weak (resp. strong) core is nomempty, then it contains a matching $M$, for which $x_{i}=|V(M)\cap V_{i}|$ gives a maximal vector $x\in\mathcal{X}$ (else, there is a matching, where everyone is weakly better off, and some players are strictly better off – so a strongly (resp. weakly) blocking coalition to the latter would also block $M$). By Corollary 21, for each $x\in\mathcal{X}$, we can check in $\mathcal{O}(2^{m}|V|^{3.5})$ time, whether every matching $M$ with $|V(M)\cap V_{i}|=x_{i}$ ($i\in[m]$) is in the strong or weak core. If the answer is negative for every $x\in\mathcal{X}$, then the core is empty; otherwise, it is not empty.
Corollary 22.
NTU-PMG-WC-Exist and NTU-PMG-SC-Exist are solvable in $\mathcal{O}(|V|^{m+2.5})$ time. Furthermore, a matching in the weak or strong core can be found in the same time, if one exists.
Optimization on the weak and strong cores can be done by similar ideas.
Theorem 23.
Given an NTU partitioned matching game with $m$ players, and a cost vector $c\in\mathbb{Q}^{E}$, we can find a minimum cost matching in the weak (resp. strong) core in $\mathcal{O}(|V|^{m+4})$ time.
Proof.
For a given vector $x\in\mathbb{Z}_{+}^{m}$ for which $x_{i}\leq|V_{i}|$ ($i\in[m]$), we can decide using Lemma 1 in $\mathcal{O}(2^{m}|V|^{2.5})$ time whether there is a coalition $\{i_{1},\dots,i_{k}\}$ and a matching $M$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M)\cap V_{i_{j}}|\geq x_{i_{j}}+1$ for every $j\in[k]$. Let $\mathcal{X}_{w}$ denote the set of vectors for which there is no such coalition. Then, $|\mathcal{X}_{w}|\leq\mathcal{O}(|V|^{m})$. Similarly, we can decide in $\mathcal{O}(2^{m}|V|^{3.5})$ time whether there is a coalition $\{i_{1},\dots,i_{k}\}$, an index $\ell\in[k]$, and a matching $M$ in the induced subgraph $G[V_{i_{1}}\cup V_{i_{2}}\cup\dots\cup V_{i_{k}}]$ such that $|V(M)\cap V_{i_{j}}|\geq x_{i_{j}}$ for every $j\in[k]$ and $|V(M)\cap V_{i_{\ell}}|\geq x_{i_{\ell}}+1$. Let $\mathcal{X}_{s}$ denote the set of vectors for which there is no such coalition. Then, $|\mathcal{X}_{s}|\leq\mathcal{O}(|V|^{m})$.
For $x\in\mathcal{X}_{w}$, let $\mathcal{M}_{x}$ denote the set of matchings in $G$ satisfying $|V(M)\cap V_{i}|\geq x_{i}$ for every $i\in[m]$. Note that $x\in\mathcal{X}_{w}$ means that every matching in $\mathcal{M}_{x}$ is in the weak core; furthermore, every matching in the weak core is in $\mathcal{M}_{x}$ for some $x\in\mathcal{X}_{w}$. Then, we can find a minimum cost matching in $\mathcal{M}_{x}$ for each $x\in\mathcal{X}_{w}$ in $\mathcal{O}(|V|^{4})$ time by Lemma 1. This, in total, gives a runtime of $\mathcal{O}(|V|^{m+4})$.
The strong core case is analogous.
@@ 4 Hardness results for partition classes of constant size
In contrast to the case where the number of players is bounded by a constant, bounding the sizes of the partition classes by a constant larger than 2 does not lead to polynomial-time solvable problems. In this section, we show hardness results of this type. We have seen in Section 2 that our problems are tractable when all partition classes have size at most 2. Here we show that having partition classes of size 3 leads to coNP-hardness for most problems.
@@ 4.1 Weak core
Our first result is the hardness of verifying if a matching belongs to the weak core.
@@
NTU-PMG-WC-Verif-3 is coNP-complete, even if the graph is bipartite.
@@
Let $I$ be an instance of x3c. We create a bipartite instance $I^{\prime}$ of the NTU-PMG-WC-Verif-3 as follows. First, we describe the vertex set $V$ of the graph.
  • For each element $i\in[3n]$ we create a vertex $a_{i}$.
@@
Finally, we describe the partition of the players. For each $j\in[m]$, we have a player $C_{j}=\{c_{j}^{1},c_{j}^{2},c_{j}^{3}\}$ and a player $D_{j}=\{d_{j}^{1},d_{j}^{2}\}$. For each $i\in\{2,\dots,3n-1\}$ we have a player $A_{i}=\{b_{i-1,i}^{i},a_{i},b_{i,i+1}^{i}\}$. In addition, we have players $A_{1}=\{a_{1},b_{1,2}^{1}\}$, and $A_{3n}=\{b_{3n-1,3n}^{3n},a_{3n}\}$. Clearly, all players have size at most 3.
We create a matching $M$ as follows. We add $c_{j}^{l}d_{j}^{l}$ for $j\in[m],l\in[2]$ and $b_{i,i+1}^{i}b_{i,i+1}^{i+1}$ for $i\in[3n-1]$. The construction is illustrated in Figure 5
To show coNP-hardness, we show that $I$ admits an exact 3-cover if and only if $M$ is not in the weak core.
@@
Let $\{S_{l_{1}},\dots S_{l_{n}}\}$ be an exact 3-cover in $I$. We show that the players $\{C_{l_{1}},\dots,C_{l_{n}},A_{1},\dots,A_{3n}\}$ form a strongly blocking coalition. As $\{S_{l_{1}},\dots S_{l_{n}}\}$ is an exact cover, there must be a perfect matching between the included $c_{j}^{l}$ and $a_{i}$ vertices. Also, we can include all edges of the form $(b_{i,i+1}^{i},b_{i,i+1}^{i+1})$ for $i=1,\dots,3n-1$. Hence, there is a perfect matching on the vertex set of these players. As each of these players has a vertex exposed in $M$ ($c_{j}^{3}$ for $C_{j}$, $a_{i}$ for $A_{i}$), it follows that $M$ is not in the weak core. ∎
@@
Suppose that $M$ is not in the weak core. This means that there is a strongly blocking coalition $\mathcal{P}$ to $M$ with a matching $M^{\prime}$. We can assume that there is a player $A_{i}\in\mathcal{P}$. Otherwise, there could be only $C_{j}$ players in $\mathcal{P}$ (the $D_{j}$ players have all vertices covered by $M$), and they induce no edges between them, a contradiction.
Observe that players $A_{1},\dots,A_{3n}$ must all be in $\mathcal{P}$, because for $A_{i}$ to strictly improve, $b_{i-1,i}^{i}$ and $b_{i,i+1}^{i}$ (if they exist) must be matched, so $b_{i-1,i}^{i-1}b_{i-1,i}^{i},b_{i,i+1}^{i}b_{i,i+1}^{i+1}\in M^{\prime}$. This implies that $A_{i-1},A_{i+1}\in\mathcal{P}$ and similarly we get that $A_{1},\dots,A_{3n}$ are all in $\mathcal{P}$.
Finally, as each player must strictly improve, every $a_{i}$ vertex is matched in $M^{\prime}$ and so they are matched to $C_{j}$ players. Whenever a $C_{j}$ player participates in the blocking coalition by matching one of its vertices to an $a_{i}$ vertex, $\{c_{j}^{1}a_{j_{1}},c_{j}^{2}a_{j_{2}},c_{j}^{3}a_{j_{3}}\}\subseteq M^{\prime}$, because $D_{j}\notin\mathcal{P}$ and $C_{j}$ must strictly improve. By combining this with the fact that each $a_{i}$ is matched, we obtain that these players must correspond to an exact 3-cover. ∎
@@
Figure 5: Illustration for Theorem 24. The figure depicts the instance $I^{\prime}$ corresponding to an x3c instance with $S_{1}=\{1,2,3\},S_{2}=\{2,4,5\}$ and $S_{3}=\{4,5,6\}$. Bold edges denote the edges of $M$.
Intuitively, it would be tempting to conjecture that the weak core is always nonempty in any NTU partitioned matching game. However, the following example (communicated to us by Zsuzsanna Jankó) shows that this is not the case.
Example 27 (Empty weak core).